# DEEP-UNFOLDING-BASED JOINT ACTIVITY AND DATA DETECTION FOR GRANT-FREE TRANSMISSION IN CELL-FREE WIRELESS COMMUNICATION SYSTEMS

### 20 Dec 2023 v1

@@ Abstract
Index Terms—  Cell-free, deep unfolding, joint activity and data detection, massive machine-type communication.
1 Introduction
@@
$\textstyle\mathcal{P}_{1}:\;\Big{\{}{\hat{\mathbf{H}},{{\hat{\mathbf{X}}}_{\text{D}}}}\Big{\}}=\mathop{\arg\min}{}_{\scriptstyle{\mathbf{H}\in\mathbb{C}^{MP\times N}}\hfill\atop{\scriptstyle{{\mathbf{X}}_{\text{D}}\in\mathcal{B}}\hfill}}\frac{1}{2}\left\|\mathbf{Y}-\mathbf{H}\left[\mathbf{X}_{\text{P}},\mathbf{X}_{\text{D}}\right]\right\|_{F}^{2}+\mu_{h}\sum_{n=1}^{N}\sum_{p=1}^{P}\left\|\mathbf{h}_{n,p}\right\|_{F}+\mu_{x}\sum_{n=1}^{N}\left\|\mathbf{x}_{\text{D},n}\right\|_{F}.$ (3)
1.1 Contributions
@@ 1.2 Relevant Prior Art
2 System Model and JAD
@@ 2.1 System Model
$\textstyle\mathbf{Y}=\sum_{n=1}^{N}\xi_{n}\mathbf{h}_{n}\bar{\mathbf{x}}_{n}^{T}+\mathbf{N}=\mathbf{H}\mathbf{X}+\mathbf{N}.$ (1)
@@ 2.2 Problem Formulation
$\textstyle\{\hat{\mathbf{H}},{{\hat{\mathbf{X}}}_{\text{D}}}\}=\mathop{\arg\max}\limits_{\scriptstyle{\mathbf{H}\in\mathbb{C}^{MP\times N}}\hfill\atop{\scriptstyle{{\mathbf{X}}_{\text{D}}\in\bar{\mathcal{Q}}}\hfill}}P\!\left(\mathbf{Y}|\mathbf{H},\mathbf{X}_{\text{D}}\right)\!P\!\left(\mathbf{H}\right)\!P\!\left(\mathbf{X}_{\text{D}}\right),$ (2)
with $P\!\left(\mathbf{Y}|\mathbf{H},\mathbf{X}_{\text{D}}\right)\propto\exp(-\left\|\mathbf{Y}-\mathbf{H}\left[\mathbf{X}_{\text{P}},\mathbf{X}_{\text{D}}\right]\right\|_{F}^{2}/N_{0})$. As in [1], we leverage the complex-valued block-Laplace model $P\!\left(\mathbf{H}\right)\propto\prod_{n=1}^{N}\prod_{p=1}^{P}\exp(-2\mu_{h}\left\|\mathbf{h}_{n,p}\right\|_{F})$ and the complex-valued Laplace model $P\!\left(\mathbf{X}_{\text{D}}\right)\propto\prod_{n=1}^{N}\exp(-2\mu_{x}\left\|\mathbf{x}_{\text{D},n}\right\|_{F})$ to take the sparsity of $\mathbf{H}$ and $\mathbf{X}_{\text{D}}$ into account [7], where $\mu_{h}$ and $\mu_{x}$ are parameters that determine the sparsity of $\mathbf{H}$ and $\mathbf{X}_{\text{D}}$, respectively. The discrete nature $\mathbf{X}_{\text{D}}$ due to the constellation $\bar{\mathcal{Q}}$ renders $\mathcal{P}_{1}$ a discrete-valued optimization problem and a naïve exhaustive search is simply infeasible. To arrive at a computationally manageable optimization problem, we relax the set $\bar{\mathcal{Q}}$ to its convex hull =$\mathcal{B}=\{\sum_{i=1}^{|\bar{\mathcal{Q}|}}\delta_{i}\mathbf{X}_{i}:\,\mathbf{X}_{i}\in\bar{\mathcal{Q}},\delta_{i}\geq 0,\forall i;\,\sum_{i=1}^{|\bar{\mathcal{Q}|}}\delta_{i}=1\}\!$ as in [7].
@@
2.3 FBS-based JAD Algorithm
With the definition $\mathbf{S}\triangleq[\mathbf{H}^{H},\mathbf{X}_{\text{D}}]^{H}\in\mathbb{C}^{(MP+R_{\text{D}})\times N}$, the FBS-based JAD algorithm starts by splitting the objective function in $\mathcal{P}_{1}$ into $f(\mathbf{S})=\textstyle\frac{1}{2}\|\mathbf{Y}-\mathbf{H}[\mathbf{X}_{\text{P}},\mathbf{X}_{\text{D}}]\|_{F}^{2}$ and $g(\mathbf{S})=\textstyle\mu_{h}\sum_{n=1}^{N}\sum_{p=1}^{P}\|\mathbf{h}_{n,p}\|_{F}+\mu_{x}\sum_{n=1}^{N}\|\mathbf{x}_{\text{D},n}\|_{F}+\mathcal{X}(\mathbf{X}_{\text{D}})$, where $\mathcal{X}\!\left(\mathbf{X}_{\text{D}}\right)\!=+\infty\,\mathbb{I}\{\exists\mathbf{X}_{\text{D}}\notin\mathcal{B}\}$ enforces the data to be within the convex set $\mathcal{B}$. FBS alternates between a gradient-descent step in the smooth function $f(\mathbf{S})$ (the forward step) and performing a proximal operation to find a solution near the minimizer of $g(\mathbf{S})$ (the backward step), where iterations are carried out until a convergence criterion is met. The forward and backward steps are as follows:
1) Forward Step: The forward step is given by $\hat{\mathbf{S}}^{k}=\mathbf{S}^{k}-\tau^{k}\nabla f\!\left(\mathbf{S}^{k}\right)\!$, where the superscript $k$ indicates the $k$th iteration, $\hat{\mathbf{S}}^{k}=[(\hat{\mathbf{H}}^{k})^{H},\hat{\mathbf{X}}_{\text{D}}^{k}]^{H}$, $\tau^{k}$ is the step size [20], and the gradient of $f\!\left(\mathbf{S}\right)\!$ with respect to $\mathbf{S}$ is given by $\nabla f(\mathbf{S})=-[\mathbf{X}^{*}\!\left(\mathbf{Y}^{T}-\mathbf{X}^{T}\mathbf{H}^{T}\right)\!,\mathbf{H}^{T}\!\left(\mathbf{Y}_{\text{D}}-\mathbf{H}\mathbf{X}_{\text{D}}\right)\!^{*}]^{T}$.
2) Backward Step: The proximal operator for $\mathbf{H}$ is $\mathbf{h}_{n,p}^{k+1}={\max\{\|\hat{\mathbf{h}}_{n,p}^{k}\|_{F}-\tau^{k}\mu_{h},0\}\hat{\mathbf{h}}_{n,p}^{k}}/{\|\hat{\mathbf{h}}_{n,p}^{k}\|_{F}}$, and the proximal operation for $\mathbf{X}_{\text{D}}$ involves derivations using the Karush-Kuhn-Tucker conditions; see [1] for the details.
3 DU-JAD: Deep-Unfolding-based joint activity and data detection
Fig. 1: Architecture of the DU-JAD algorithm.
@@ 3.1 FBS Module
$\hat{\mathbf{H}}^{k}=\mathbf{H}^{k}+\tau_{h}^{k}\!\left(\mathbf{Y}-\mathbf{H}^{k}\mathbf{X}^{k}\right)\!(\mathbf{X}^{k})^{H}+\eta_{h}^{k}\mathbf{D}_{h}^{k-1},$ (4)
$\hat{\mathbf{X}}_{\text{D}}^{k}=\mathbf{X}_{\text{D}}^{k}+\tau_{x}^{k}\!\left(\mathbf{Y}_{\text{D}}-\mathbf{H}^{k}\mathbf{X}_{\text{D}}^{k}\right)\!^{H}\mathbf{H}^{k}+\eta_{x}^{k}\mathbf{D}_{x}^{k-1},$
where $\mathbf{D}_{h}^{k}=\tau_{h}^{k}\!\left(\mathbf{Y}-\mathbf{H}^{k}\mathbf{X}^{k}\right)\!(\mathbf{X}^{k})^{H}+\eta_{h}^{k}\mathbf{D}_{h}^{k-1}$ and $\mathbf{D}_{x}^{k}=\tau_{x}^{k}\!\left(\mathbf{Y}_{\text{D}}-\mathbf{H}^{k}\mathbf{X}_{\text{D}}^{k}\right)\!^{H}\mathbf{H}^{k}+\eta_{x}^{k}\mathbf{D}_{x}^{k-1}$ with $\mathbf{D}_{h}^{0}=\mathbf{0}$ and $\mathbf{D}_{x}^{0}=~{}\mathbf{0}$.
@@
We start by rewriting the $n$th UE’s estimated data vector $\hat{\mathbf{x}}_{\text{D},n}^{k}$ in the $k$th iteration as $\hat{\mathbf{x}}_{\text{D},n}^{k}=\mathbf{x}_{\text{D},n}+\mathbf{e}^{k}$, where $\mathbf{x}_{\text{D},n}$ is the unknown data vector and $\mathbf{e}^{k}$ is the estimation error in iteration $k$. By assuming $\mathbf{e}^{k}\sim\mathcal{CN}\left(\mathbf{0},N_{\text{e}}^{k}\mathbf{I}_{R_{\text{D}}}\right)$, the PME of $\mathbf{x}_{\text{D},n}$ is given by
$\mathbf{x}_{\text{D},n}^{k}=\frac{\sum_{\mathbf{x}_{\text{D},n}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x}_{\text{D},n})\,{P_{a}}/{|\mathcal{Q}|^{R_{\text{D}}}}\mathbf{x}_{\text{D},n}}{\sum_{\mathbf{x}_{\text{D},n}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x}_{\text{D},n})\,{P_{a}}/{|\mathcal{Q}|^{R_{\text{D}}}}+P_{\text{PME}}^{n,k}(\mathbf{0})(1-P_{a})},$ (5)
@@
where $P_{\text{PME}}^{n,k}(\mathbf{x})\triangleq\exp(-{\|\mathbf{x}-\hat{\mathbf{x}}_{\text{D},n}^{k}\|_{F}^{2}}/{N_{\text{e}}^{k}})$, and $P_{a}$ represents the user activity probability following $P(\mathbf{X}_{\text{D}})$. Evaluating the PME expression of $\mathbf{x}_{\text{D},n}$ incurs high complexity and prone to numerical stability issues. Consequently, we employ an approximate PME as $\mathbf{x}_{\text{D},n}^{k}=\alpha_{n}^{k}\,\check{\mathbf{x}}_{\text{D},n}^{k}$. Here, the scaling factor $\alpha_{n}^{k}\in[0,1]$ can be expressed as
$\alpha_{n}^{k}=\min\big{\{}\frac{1}{\|\hat{\mathbf{x}}_{\text{D},n}^{k}\|_{1}}\max\{\lambda^{k}\|\hat{\mathbf{x}}_{\text{D},n}^{k}\|_{1}-\nu^{k},0\},1\big{\}},$ (6)
@@
where $\lambda^{k}$ and $\nu^{k}$ are trainable parameters. Besides, $\check{\mathbf{x}}_{\text{D},n}^{k}\triangleq\\ {\sum_{\mathbf{x}_{\text{D},n}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x}_{\text{D},n})\,\mathbf{x}_{\text{D},n}}/{\sum_{\mathbf{x}_{\text{D},n}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x}_{\text{D},n})}$ with the estimation error variance $N_{\text{e}}^{k}$ to be trained by DU.
3.2 AUD and DD Modules
@@
To improve AUD, we propose a soft-output module that delivers the probability of each user being active instead of carrying out hard-output decisions. To this end, we propose to use a sigmoid within the AUD module, which leverages the estimated $\mathbf{S}^{K+1}$ to output the likelihood of the user activity status as $L_{n}=1/(1+1/\exp(\omega_{h}\|\mathbf{h}_{n}^{K+1}\|_{F}^{2}+\omega_{x}\|\mathbf{x}_{\text{D},n}^{K+1}\|_{F}^{2}-T_{\text{th}}))$, with trainable parameters $\omega_{h}$, $\omega_{x}$ and $T_{\text{th}}$. From the calculated activity probability, the user’s active state can be detected by comparing it to a threshold $\bar{L}\in[0,1]$, that is $\hat{\xi}_{n}=\mathbb{I}\left\{L_{n}>\bar{L}\right\}$. As for the DD module, we can detect the data matrix as $\hat{\mathbf{X}}_{\text{D}}=\operatorname{diag}\{\hat{\xi}_{1},\ldots,\hat{\xi}_{N}\}\tilde{\mathbf{X}}_{\text{D}}$, where $\tilde{\mathbf{X}}_{\text{D}}=\arg\min_{\mathbf{X}\in\mathcal{Q}^{N\times R_{\text{D}}}}\|\mathbf{X}-\mathbf{X}_{\text{D}}^{K+1}\|_{F}^{2}$.
3.3 Training Procedure

### 28 Feb 2024 v2

# Deep-Unfolded Joint Activity and Data Detection for Grant-Free Transmission in Cell-Free Systems

@@ Abstract
Index Terms:
Cell-free, deep unfolding, joint activity and data detection, massive machine-type communication.
I Introduction
@@
$\mathcal{P}_{1}:\;\Big{\{}{\hat{\mathbf{H}},{{\hat{\mathbf{X}}}_{\text{D}}}}\Big{\}}=\mathop{\arg\min}_{\scriptstyle{\mathbf{H}\in\mathbb{C}^{MP\times N}}\hfill\atop{\scriptstyle{{\mathbf{X}}_{\text{D}}\in\mathcal{B}}\hfill}}\frac{1}{2}\left\|\mathbf{Y}-\mathbf{H}\left[\mathbf{X}_{\text{P}},\mathbf{X}_{\text{D}}\right]\right\|_{F}^{2}+\mu_{h}\sum_{n=1}^{N}\sum_{p=1}^{P}\left\|\mathbf{h}_{n,p}\right\|_{F}+\mu_{x}\sum_{n=1}^{N}\left\|\mathbf{x}_{\text{D},n}\right\|_{F}.$ (7)
I-A Contributions
@@ I-B Relevant Prior Art
I-C Notation
Matrices are denoted by uppercase boldface letters, column vectors by lowercase boldface letters, and sets by uppercase calligraphic letters. The entry at the $m$th row and $n$th column of matrix $\mathbf{A}$ is $\mathbf{A}(m,n)$ ;the $m$th row vector of matrix $\mathbf{A}$ is $\mathbf{A}(m,:)$. The superscripts $(\cdot)^{*}$, $(\cdot)^{T}$, and $(\cdot)^{H}$ denote conjugate, transpose, and conjugate transpose, respectively; $\mathbf{I}_{N}$ is the $N\times N$ identity matrix. A diagonal matrix with entries $\{x_{1},\ldots,x_{N}\}$ on the main diagonal is $\operatorname{diag}\{x_{1},\ldots,x_{N}\}$. The indicator function $\mathbb{I}\{\cdot\}$ is one if its condition is true and zero otherwise. The $\ell_{1}$ norm and Frobenius norm are $\|\cdot\|_{1}$ and $\|\cdot\|_{F}$, respectively. The cardinality of the set $\mathcal{Q}$ is $|\mathcal{Q}|$.
II System Model and JAD
@@ II-A System Model
$\mathbf{Y}=\sum_{n=1}^{N}\xi_{n}\mathbf{h}_{n}\bar{\mathbf{x}}_{n}^{T}+\mathbf{N}=\mathbf{H}\mathbf{X}+\mathbf{N}.$ (1)
@@ II-B Problem Formulation
$\{\hat{\mathbf{H}},{{\hat{\mathbf{X}}}_{\text{D}}}\}=\mathop{\arg\max}\limits_{\scriptstyle{\mathbf{H}\in\mathbb{C}^{MP\times N}}\hfill\atop{\scriptstyle{{\mathbf{X}}_{\text{D}}\in\bar{\mathcal{Q}}}\hfill}}P\!\left(\mathbf{Y}|\mathbf{H},\mathbf{X}_{\text{D}}\right)\!P\!\left(\mathbf{H}\right)\!P\!\left(\mathbf{X}_{\text{D}}\right),$ (2)
with
$P\!\left(\mathbf{Y}|\mathbf{H},\mathbf{X}_{\text{D}}\right)\propto\exp\left(-\frac{\left\|\mathbf{Y}-\mathbf{H}\left[\mathbf{X}_{\text{P}},\mathbf{X}_{\text{D}}\right]\right\|_{F}^{2}}{N_{0}}\right)\!.$ (3)
As in [1] and [7], we leverage the following complex-valued block-Laplace model and complex-valued Laplace model to take the sparsity of $\mathbf{H}$ and $\mathbf{X}_{\text{D}}$ into account:
$P\!\left(\mathbf{H}\right)\propto\prod_{n=1}^{N}\prod_{p=1}^{P}\exp\left(-2\mu_{h}\left\|\mathbf{h}_{n,p}\right\|_{F}\right),$ (4)
$P\!\left(\mathbf{X}_{\text{D}}\right)\propto\prod_{n=1}^{N}\exp\left(-2\mu_{x}\left\|\mathbf{x}_{\text{D},n}\right\|_{F}\right),$ (5)
where $\mu_{h}$ and $\mu_{x}$ are parameters that determine the sparsity of $\mathbf{H}$ and $\mathbf{X}_{\text{D}}$, respectively. The discrete nature $\mathbf{X}_{\text{D}}$ due to the constellation $\bar{\mathcal{Q}}$ renders $\mathcal{P}_{1}$ a discrete-valued optimization problem and a naïve exhaustive search is simply infeasible. To arrive at a computationally manageable optimization problem, we relax the set $\bar{\mathcal{Q}}$ to its convex hull [7]:
=$\mathcal{B}=\left\{\sum_{i=1}^{|\bar{\mathcal{Q}|}}\delta_{i}\mathbf{X}_{i}:\,\mathbf{X}_{i}\in\bar{\mathcal{Q}},\delta_{i}\geq 0,\forall i;\,\sum_{i=1}^{|\bar{\mathcal{Q}|}}\delta_{i}=1\right\}\!.$ (6)
@@
II-C FBS-based JAD Algorithm
With the definition $\mathbf{S}\triangleq[\mathbf{H}^{H},\mathbf{X}_{\text{D}}]^{H}\in\mathbb{C}^{(MP+R_{\text{D}})\times N}$, the FBS-based JAD algorithm starts by splitting the objective function in $\mathcal{P}_{1}$ into
$f(\mathbf{S})=\frac{1}{2}\left\|\mathbf{Y}-\mathbf{H}\left[\mathbf{X}_{\text{P}},\mathbf{X}_{\text{D}}\right]\right\|_{F}^{2},$ (8)
$g(\mathbf{S})=\mu_{h}\sum_{n=1}^{N}\sum_{p=1}^{P}\left\|\mathbf{h}_{n,p}\right\|_{F}+\mu_{x}\sum_{n=1}^{N}\left\|\mathbf{x}_{\text{D},n}\right\|_{F}+\mathcal{X}(\mathbf{X}_{\text{D}}),$ (9)
where $\mathcal{X}\!\left(\mathbf{X}_{\text{D}}\right)\!=+\infty\,\mathbb{I}\{\exists\mathbf{X}_{\text{D}}\notin\mathcal{B}\}$ enforces the data to be within the convex set $\mathcal{B}$. FBS alternates between a gradient-descent step in the smooth function $f(\mathbf{S})$ (the forward step) and performing a proximal operation to find a solution near the minimizer of $g(\mathbf{S})$ (the backward step), where iterations are carried out until a convergence criterion is met. The forward and backward steps are as follows:
1) Forward Step: The forward step is given by
$\hat{\mathbf{S}}^{k}=\mathbf{S}^{k}-\tau^{k}\nabla f\!\left(\mathbf{S}^{k}\right)\!,$ (10)
where the superscript $k$ indicates the $k$th iteration, $\hat{\mathbf{S}}^{k}=[(\hat{\mathbf{H}}^{k})^{H},\hat{\mathbf{X}}_{\text{D}}^{k}]^{H}$, $\tau^{k}$ is the step size [20], and the gradient of $f\!\left(\mathbf{S}\right)\!$ with respect to $\mathbf{S}$ is given by
$\nabla f(\mathbf{S})=-\left[\mathbf{X}^{*}\!\left(\mathbf{Y}^{T}-\mathbf{X}^{T}\mathbf{H}^{T}\right)\!,\mathbf{H}^{T}\!\left(\mathbf{Y}_{\text{D}}-\mathbf{H}\mathbf{X}_{\text{D}}\right)\!^{*}\right]^{T}.$ (11)
2) Backward Step: The proximal operator for $\mathbf{H}$ is
$\mathbf{h}_{n,p}^{k+1}=\frac{\max\left\{\left\|\hat{\mathbf{h}}_{n,p}^{k}\right\|_{F}-\tau^{k}\mu_{h},0\right\}\hat{\mathbf{h}}_{n,p}^{k}}{\left\|\hat{\mathbf{h}}_{n,p}^{k}\right\|_{F}},$ (12)
and the proximal operation for $\mathbf{X}_{\text{D}}$ involves derivations using the Karush-Kuhn-Tucker conditions; see [1] for the details.
III DU-JAD: Deep-Unfolding-based joint activity and data detection
Figure 1: Architecture of the DU-JAD algorithm.
@@ III-A FBS Module
$\hat{\mathbf{H}}^{k}=\mathbf{H}^{k}+\tau_{h}^{k}\!\left(\mathbf{Y}-\mathbf{H}^{k}\mathbf{X}^{k}\right)\!\left(\mathbf{X}^{k}\right)^{H}+\eta_{h}^{k}\mathbf{D}_{h}^{k-1},$ (13)
$\hat{\mathbf{X}}_{\text{D}}^{k}=\mathbf{X}_{\text{D}}^{k}+\tau_{x}^{k}\!\left(\mathbf{Y}_{\text{D}}-\mathbf{H}^{k}\mathbf{X}_{\text{D}}^{k}\right)\!^{H}\mathbf{H}^{k}+\eta_{x}^{k}\mathbf{D}_{x}^{k-1},$ (14)
where
$\mathbf{D}_{h}^{k}=\tau_{h}^{k}\!\left(\mathbf{Y}-\mathbf{H}^{k}\mathbf{X}^{k}\right)\!\left(\mathbf{X}^{k}\right)^{H}+\eta_{h}^{k}\mathbf{D}_{h}^{k-1},$ (15)
$\mathbf{D}_{x}^{k}=\tau_{x}^{k}\!\left(\mathbf{Y}_{\text{D}}-\mathbf{H}^{k}\mathbf{X}_{\text{D}}^{k}\right)\!^{H}\mathbf{H}^{k}+\eta_{x}^{k}\mathbf{D}_{x}^{k-1},$ (16)
with $\mathbf{D}_{h}^{0}=\mathbf{0}$ and $\mathbf{D}_{x}^{0}=~{}\mathbf{0}$.
@@
We start by rewriting the $n$th UE’s estimated data vector $\hat{\mathbf{x}}_{\text{D},n}^{k}$ in the $k$th iteration as $\hat{\mathbf{x}}_{\text{D},n}^{k}=\mathbf{x}_{\text{D},n}+\mathbf{e}^{k}$, where $\mathbf{x}_{\text{D},n}$ is the unknown data vector and $\mathbf{e}^{k}$ is the estimation error in iteration $k$. By assuming $\mathbf{e}^{k}$ follows a complex Gaussian distribution with mean vector $\mathbf{0}$ and covariance matrix $N_{\text{e}}^{k}\mathbf{I}_{R_{\text{D}}}$, the PME of $\mathbf{x}_{\text{D},n}$ is given by
$\mathbf{x}_{\text{D},n}^{k}=\frac{\sum_{\mathbf{x}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x})\,{P_{a}}/{|\mathcal{Q}^{R_{\text{D}}}|}\,\mathbf{x}}{\sum_{\mathbf{x}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x})\,{P_{a}}/{|\mathcal{Q}^{R_{\text{D}}}}|+P_{\text{PME}}^{n,k}(\mathbf{0})(1-P_{a})},$ (17)
where $P_{a}$ represents the user activity probability following $P(\mathbf{X}_{\text{D}})$, and the function $P_{\text{PME}}^{n,k}(\mathbf{x})$ is defined as
$P_{\text{PME}}^{n,k}(\mathbf{x})\triangleq\exp\left(-\frac{\left\|\mathbf{x}-\hat{\mathbf{x}}_{\text{D},n}^{k}\right\|_{F}^{2}}{N_{\text{e}}^{k}}\right).$ (18)
@@
Evaluating the PME expression of $\mathbf{x}_{\text{D},n}$ incurs high complexity and prone to numerical stability issues. Consequently, we employ an approximate PME as $\mathbf{x}_{\text{D},n}^{k}=\alpha_{n}^{k}\,\check{\mathbf{x}}_{\text{D},n}^{k}$. Here, the scaling factor $\alpha_{n}^{k}\in[0,1]$ can be expressed as
$\alpha_{n}^{k}=\min\left\{\frac{1}{\left\|\hat{\mathbf{x}}_{\text{D},n}^{k}\right\|_{1}}\max\{\lambda^{k}\|\hat{\mathbf{x}}_{\text{D},n}^{k}\|_{1}-\nu^{k},0\},1\right\},$ (19)
@@
where $\lambda^{k}$ and $\nu^{k}$ are trainable parameters. Besides, the vector $\check{\mathbf{x}}_{\text{D},n}^{k}$ is defined as
$\check{\mathbf{x}}_{\text{D},n}^{k}\triangleq\frac{\sum_{\mathbf{x}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x})\,\mathbf{x}}{\sum_{\mathbf{x}\in\mathcal{Q}^{R_{\text{D}}}}P_{\text{PME}}^{n,k}(\mathbf{x})},$ (20)
with the estimation error variance $N_{\text{e}}^{k}$ to be trained by DU.
III-B AUD and DD Modules
@@
To improve AUD, we propose a soft-output module that delivers the probability of each user being active instead of carrying out hard-output decisions. To this end, we propose to use a sigmoid within the AUD module, which leverages the estimated $\mathbf{S}^{K+1}$ to output the likelihood of the user activity status as
$L_{n}=\frac{1}{1+1/\exp\left(\omega_{h}\left\|\mathbf{h}_{n}^{K+1}\right\|_{F}^{2}+\omega_{x}\left\|\mathbf{x}_{\text{D},n}^{K+1}\right\|_{F}^{2}-T_{\text{th}}\right)},$ (21)
with trainable parameters $\omega_{h}$, $\omega_{x}$ and $T_{\text{th}}$. From the calculated activity probability, the user’s active state can be detected by comparing it to a threshold $\bar{L}\in[0,1]$, i.e.,
$\hat{\xi}_{n}=\mathbb{I}\left\{L_{n}>\bar{L}\right\}.$ (22)
As for the DD module, we detect the data matrix as
$\hat{\mathbf{X}}_{\text{D}}=\operatorname{diag}\{\hat{\xi}_{1},\ldots,\hat{\xi}_{N}\}\tilde{\mathbf{X}}_{\text{D}},$ (23)
where $\tilde{\mathbf{X}}_{\text{D}}=\arg\min_{\mathbf{X}\in\mathcal{Q}^{N\times R_{\text{D}}}}\|\mathbf{X}-\mathbf{X}_{\text{D}}^{K+1}\|_{F}^{2}$.
III-C Training Procedure